A gate is to be constructed with the following description: 46 1/4 in. wide. 10 boards, 3 1/2 in. wide, placed vertically and equally spaced Dete

Algebra ->  Customizable Word Problem Solvers  -> Misc -> SOLUTION: Ref. Question 1045604: A gate is to be constructed with the following description: 46 1/4 in. wide. 10 boards, 3 1/2 in. wide, placed vertically and equally spaced Dete      Log On

Ad: Over 600 Algebra Word Problems at edhelper.com


   



Question 1045894: Ref. Question 1045604:
A gate is to be constructed with the following description:
46 1/4 in. wide.
10 boards, 3 1/2 in. wide, placed vertically and equally spaced
Determine space between boards.

Answer by Alan3354(69443)

A gate is to be constructed with the following description:
46 1/4 in. wide.
10 boards, 3 1/2 in. wide, placed vertically and equally spaced
Determine space between boards.
--------------
10 boards & 9 spaces between them.
10*3.5 = 35 inches.
46.25 - 35 = 11.25 inches
11.25/9 = 1.25" spaces